OpenStudy (anonymous):

Algebra 1 help please! I'll give a medal to anyone who helps me. :) * * * Question: An expression is shown below: 3x3y + 12xy - 9x2y - 36y Part A: Rewrite the expression so that the GCF is factored completely. Show the steps of your work. Part B: Rewrite the expression completely factored. Show the steps of your work. Part C: If the two middle terms were switched so that the expression became 3x3y - 9x2y + 12xy - 36y, would the factored expression no longer be equivalent to your answer in part B? Explain your reasoning.

ganeshie8 (ganeshie8):

\(\large 3x^3y + 12xy - 9x^2y - 36y\) GCF of numbers = 3 GCF of y terms = y pull them out, what do u get ?

ganeshie8 (ganeshie8):

\(\large 3x^3y + 12xy - 9x^2y - 36y\) after factoring the 3 : \(\large 3(x^3y + 4xy - 3x^2y - 12y)\) after factoring the y : \(\large 3y(x^3 + 4x - 3x^2 - 12)\)

ganeshie8 (ganeshie8):

^^ thats ur part A see if that makes sense...

OpenStudy (anonymous):

I'm confused about factoring y? how did you get that? Everything else makes sense. :)

ganeshie8 (ganeshie8):

good question :)

ganeshie8 (ganeshie8):

look at the original given expression

ganeshie8 (ganeshie8):

It has 4 terms, and each term has a "y" in it right ?

ganeshie8 (ganeshie8):

Here is the given expression : \(\large 3x^3\color{red}{y} + 12x\color{red}{y} - 9x^2\color{red}{y} - 36\color{red}{y} \)

ganeshie8 (ganeshie8):

I have just pulled that "y" out as it is common in every term..

OpenStudy (anonymous):

ohhh now it makes sense.

ganeshie8 (ganeshie8):

good :) see if u can do the part B now

ganeshie8 (ganeshie8):

for part B, u need to take the Part A answer and start from there..

OpenStudy (anonymous):

okay :)

ganeshie8 (ganeshie8):

part B : \(\large 3y(x^3 + 4x - 3x^2 - 12) \) try to factor the stuff inside parenthesis

ganeshie8 (ganeshie8):

factor first two terms, and last terms separately

ganeshie8 (ganeshie8):

whats the GCF of first two terms ? \(\large 3y(\color{red}{x^3 + 4x} - 3x^2 - 12) \)

ganeshie8 (ganeshie8):

they both have a "x" in them right ?

OpenStudy (anonymous):

Yeah, do I have to take them out?

ganeshie8 (ganeshie8):

yup :) \(\large 3y(\color{red}{x^3 + 4x} - 3x^2 - 12) \) Factoring out GCF from first two terms gives : \(\large 3y\left(\color{red}{x(x^2 + 4)} - 3x^2 - 12\right) \)

ganeshie8 (ganeshie8):

Similarly, factor the GCF from last terms, you wud get : \(\large 3y(\color{red}{x^3 + 4x} - 3x^2 - 12) \) Factoring out GCF from first two terms gives : \(\large 3y\left(\color{red}{x(x^2 + 4)} - 3x^2 - 12\right) \) Factoring the GCF from last two terms gives : \(\large 3y\left(\color{red}{x(x^2 + 4)} - 3(x^2 + 4)\right) \)

ganeshie8 (ganeshie8):

fine, so far eh ?

OpenStudy (anonymous):

Yes. :) So now I would have to factor which expression?

ganeshie8 (ganeshie8):

zoom out a bit and stare at the expression again

ganeshie8 (ganeshie8):

do u see (x^2+4) common inside the parenthesis ?

ganeshie8 (ganeshie8):

\(\large 3y(\color{red}{x^3 + 4x} - 3x^2 - 12) \) Factoring out GCF from first two terms gives : \(\large 3y\left(\color{red}{x(x^2 + 4)} - 3x^2 - 12\right) \) Factoring the GCF from last two terms gives : \(\large 3y\left(\color{red}{x(x^2 + 4)} - 3(x^2 + 4)\right) \) factoring out the (x^2+4) from the terms inside parenthesis gives : \(\large 3y(x^2 + 4)(x-3) \)

ganeshie8 (ganeshie8):

we're done with part B !

ganeshie8 (ganeshie8):

let me knw if smthng doesnt make sense..

OpenStudy (anonymous):

Yay! Thanks so much for your help. :)

ganeshie8 (ganeshie8):

np :) what do u think about part C ?

OpenStudy (anonymous):

I think I can do it. :) aha

ganeshie8 (ganeshie8):

Very good :)